# Operationalizing content moderation “accuracy” in the Digital Services Act

### v3

@@ C.1. Stratification with different classifiers
We study three models: a unigram logistic regression model, Roberta (Liu et al., 2019), and Distilbert. Roberta (Liu et al., 2019) is a transformer-based language model which is pretrained on a large corpus of English data. Distilbert is smaller version of Roberta with the same model architecture but also 40% less parameters. We find both pretrained models to have strong performance for civilcomments after finetuning. We use typical hyperparameters for small datasets (Mosbach et al., 2021), and make two adjustments that we found to be critical: increasing the batch size to 32 (we hypothesize that since the labels can be noisy, the gradients are unstable), and subsampling the negative examples so that the training data is balanced (the resulting training data has 11k examples).
@@ Results
Figure 2. Number of samples required to estimate the prevalence of toxicity within 20% in civilcomments, with different prevalences, with different number of strata. Binning is based on predicted scores from a finetuned Roberta classifier. All pilot results are averaged over 100 trials. The numbers associated with the pilot is the total number of pilot samples across all bins. For small prevalences, more pilot samples are better, because more samples estimate the standard deviations within each bin better. In contrast, more bins fares worse because each bin has less pilot samples.
C.2. Stratification for rarer prevalences
For each example, civilcomments has a real valued label of toxicity corresponding to the proportion of annotators that labelled it as toxic. When converting the labels to binary, a threshold of 0.5 yields a prevalence of 5.9%, which is what we report on in §5. Here we present results for different prevalences by varying the threshold among $\{0.48,0.83,0.95\}$. We vary the number of total pilot samples and bins, and by choosing appropriate powers of two, each bin will always have an integer number of pilot samples.
Refer to Figure 2. In general, stratified sampling provides significant efficiency over random sampling (comparing to the figures in Table 2). The number of pilot samples collected matters for both very small and very large prevalences. In the case of high prevalence, pilot samples outnumber the optimal allocation so there will be wasted samples. In the case of low prevalence, additional pilot samples help estimate the standard deviation of the strata correctly for optimal allocation. Finally, increasing the number of strata does not help with lower prevalences, as less samples are available for each bin to estimate the standard deviation with.
@@ Appendix D Additional annotation details
We label a comment as a personal attack if it fits any of criteria defined in (Habernal et al., 2018, Table 2) including: vulgar insult, illiteracy insult, ridicule or sarcasm, condescension, accusation of stupidity, accusation of ignorance, and accusation of argumentative fallacies. We found most personal attacks to be immediately recognizable because it included insults or accusations. For borderline cases, such as rude comments, we labelled it a personal attack if it didn’t engage with the parent comment. In some other cases, viewing the original comment thread of Reddit was helpful. When the comment was part of a “flame war”, we would annotate it as a personal attack. All annotation was conducted by trusted and compensated annotator, and the rate of annotation was about 600 examples per hour. In a small interannotator agreement study with the first author, the agreement was substantial, with a Cohen’s $\kappa$ of 0.69 (out of 100 examples, both annotators agreed on 96).

### v5

@@ C.1 Stratification with different classifiers
We study three models: a unigram logistic regression model, Roberta (Liu et al. 2019), and Distilbert. Roberta (Liu et al. 2019) is a transformer-based language model which is pretrained on a large corpus of English data. Distilbert is smaller version of Roberta with the same model architecture but also 40% less parameters. We find both pretrained models to have strong performance for civilcomments after finetuning. We use typical hyperparameters for small datasets (Mosbach, Andriushchenko, and Klakow 2021), and make two adjustments that we found to be critical: increasing the batch size to 32 (we hypothesize that since the labels can be noisy, the gradients are unstable), and subsampling the negative examples so that the training data is balanced (the resulting training data has 11k examples).
@@ Appendix D Additional annotation details
We label a comment as a personal attack if it fits any of criteria defined in (Habernal et al. 2018, Table 2) including: vulgar insult, illiteracy insult, ridicule or sarcasm, condescension, accusation of stupidity, accusation of ignorance, and accusation of argumentative fallacies. We found most personal attacks to be immediately recognizable because it included insults or accusations. For borderline cases, such as rude comments, we labelled it a personal attack if it didn’t engage with the parent comment. In some other cases, viewing the original comment thread of Reddit was helpful. When the comment was part of a “flame war”, we would annotate it as a personal attack. All annotation was conducted by trusted and compensated annotator, and the rate of annotation was about 600 examples per hour. In a small interannotator agreement study with the first author, the agreement was substantial, with a Cohen’s $\kappa$ of 0.69 (out of 100 examples, both annotators agreed on 96).